OpenStudy (anonymous):

If y=4cos(2x) then d^2y/dx^2=? A.-8cos2x B.16cos2x C.-32cos2x D.-4cos2x E.-16cos2x

OpenStudy (anonymous):

You should be able to do that one with the help I gave you on the last one, just chain rule and product rule. I'll tell you if you get it right though. :P

OpenStudy (anonymous):

But it's squared :O is it still the same ?@Mikogekz

OpenStudy (anonymous):

d^2y/dx^2 is just the second derivative, all you need to do is take the derivative of the derivative.

OpenStudy (anonymous):

My first derivative is 4cos2x-88sin2x

OpenStudy (anonymous):

Only one 8 @Mikogekz

OpenStudy (anonymous):

-8sin2x is the first derivative.

OpenStudy (anonymous):

\[y=4\cos(2x)\] \[\frac{dy}{dx}=\frac{d(4\cos(2x))}{dx}\] \[=4\frac{d(\cos2x)}{dx}\] \[=4[(-\sin2x)(2)]=-8\sin(2x)\] \[\frac{dy}{dx}=-8\sin(2x)\] \[\frac{d^{2}y}{dx^{2}}=-8\frac{d(\sin2x)}{dx}\] \[=-8(\cos(2x)(2))=-16\cos(2x)\]

OpenStudy (anonymous):

nope its E...

OpenStudy (anonymous):

Forgot it was negative oops yup it's E

OpenStudy (anonymous):

Thanks again @miko